# Interpreting Graph Inference with Skyline Explanations

### v3

@@ III-A Skyline Explanatory Query
That is, ${\mathcal{G}}_{\zeta}$ is a Pareto set [24] of the interpretable space $\zeta$.
@@ III-C Computational Complexity
Proof sketch: Given an $\mathsf{SXQ}$} = $(G,{\mathcal{M}},v_{t},\Phi)$ and a set of subgraphs $\mathcal{G}$} of $G$, we provide a procedure, denoted as $\mathsf{Verify}$, that correctly determines if $\mathcal{G}$} is an explanation. The algorithm checks, for each $G_{s}\in{\mathcal{G}}$, if $G_{s}$ is a factual or a counterfactual explanation of ${\mathcal{M}}(v_{t},G)$. It has been verified that this process can be performed by invoking a polynomial time inference process of $\mathcal{M}$} [43, 37] for $v_{t}$ over $G_{s}$ (for testing factual explanation) and $G\setminus G_{s}$ (for testing counterfactual explanations), respectively. $\Box$
@@
Proof sketch: It suffices to show that ${\mathsf{EVAL}}({\mathsf{SXQ}}^{k})$ is $\mathsf{NP}$-hard for an interpretable space $|\zeta|$ = $f(|G|)$ for some polynomial function $f$, and $|\Phi|=3$. We show this by constructing a polynomial-time reduction from $k$-representative skyline selection problem ($k$-RSP). Given a set $S$ of $d$-dimensional data points, $k$-RSP is to compute a subset $S^{*}$ of $S$, such that (a) $S^{*}$ is a set of skyline points, $|S^{*}|=k$, and (b) $S^{*}$ maximizes the dominance score. $k$-RSP is $\mathsf{NP}$-hard even for $3$-dimensional data points [22].
@@
A straightforward evaluation of ${\mathsf{SXQ}}^{k}$ may explore the $L$-hop subgraph $G^{L}(\{v_{t}\})$ and initializes the interpretable space $\zeta$ as all connected subgraphs in $G^{L}(\{v_{t}\})$, by invoking subgraph enumeration algorithms [44, 45]. It then enumerates $n$ size-$k$ subsets of $\zeta$ and finds an optimal $k$-explanation. Although this correctly finds optimal explanations, it is not practical for large $G$, as $n$ alone can be $2^{deg^{L}}$ and the Pareto sets inspection can take $O(\binom{n}{k})$ time. We thus consider “approximate query processing” scheme for ${\mathsf{SXQ}}^{k}$, and present efficient algorithms that do not require such enumeration.
@@ IV-A Approximating Skyline Explanations
In other words, a $(\zeta,\epsilon)$-explanation ${\mathcal{G}}_{\epsilon}$ approximates a $k$-explanation ${\mathcal{G}}_{\zeta}$ as its answer in the interpretable space $\zeta$. Indeed, (1) ${\mathcal{G}}_{\epsilon}$ has a bounded number $k$ of explanatory subgraphs as ${\mathcal{G}}_{\zeta}$; (2) ${\mathcal{G}}_{\epsilon}$ is, by definition, an $\epsilon$-Pareto set [46] of $\zeta$. In multi-objective decision making, $\epsilon$-Pareto sets have been verified as a class of cost-effective, size-bounded approximation for Pareto optimal solutions.
@@
Update Dominance Relations. Procedure updateSX maintains a set of skyline explanations ${\mathcal{G}}_{s}$ (not shown) in terms of $(1+\epsilon)$-dominance. The set ${\mathcal{G}}_{s}$ is efficiently derived from the bitvectors $B(s)$ of the states in $\zeta^{\prime}$. The latter compactly encodes a lattice structure of $(1+\epsilon)$ dominance as a directed acyclic graph; and ${\mathcal{G}}_{s}$ refers to the states with no “parent” in the lattice, i.e., having an explanatory subgraph that is not $(1+\epsilon)$-dominated by any others so far. Details are in [47].
@@
Proof sketch: We show the above result with a reduction to the multi-objective shortest path problem ($\mathsf{MOSP}$[48]. Given an edge-weighted graph $G_{w}$, where each edge carries a $d$-dimensional attribute vector $e_{w}.c$, it computes a Pareto set of paths from a start node $u$. The cost of a path $\rho_{w}$ in $G_{w}$ is defined as $\rho_{w}.c$ = $\sum_{e_{w}\in\rho_{w}}$ $e_{w}.c$. The dominance relation between two paths is determined by the dominance relation of their cost vector. Our reduction (1) constructs $G_{w}$ as the running graph $\zeta^{\prime}$ with $n$ verified states and transitions; and (2) for each edge $(s,s^{\prime})$, sets an edge weight as $e_{w}$ = $\Phi(s)-\Phi(s^{\prime})$. Given a solution $\Pi_{w}$ of the above instance of $\mathsf{MOSP}$, for each path $\rho_{w}\in\Pi$, we set a corresponding path $\rho$ in $\zeta^{\prime}$ that ends at a state $\rho_{s}$ , and adds it into ${\mathcal{G}}_{\epsilon}$. We can verify that $\Pi_{w}$ is an $\epsilon$-Pareto set of paths $\Pi_{w}$ in $G_{w}$, if and only if ${\mathcal{G}}_{\epsilon}$ is an $(\zeta^{\prime},\epsilon)$-explanation of $\zeta^{\prime}$. We show that $\mathsf{ASX}$-$\mathsf{OP}$ performs a simpler process of the algorithm in [48], which ensures to generate ${\mathcal{G}}_{\epsilon}$ as a $(\zeta^{\prime},\epsilon)$-explanation. $\Box$
@@
Proof sketch: Consider procedure updateSX upon the arrival, at any time, of a new verified candidate $G_{s}$. (1) The above results clearly hold when $|\zeta^{\prime}|\leq k$ or $|{\mathcal{G}}_{\epsilon}|\leq k$, as ${\mathcal{G}}_{\epsilon}$ is the only $(\zeta^{\prime},\epsilon)$-explanation so far. (2) When $|{\mathcal{G}}_{\epsilon}|=k$, we reduce the approximate evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the online MAX $k$-SET problem [49]. The problem maintains a size-$k$ set cover with maximized weights. We show that updateSX adopts a greedy replacement policy by replacing a candidate in ${\mathcal{G}}_{\epsilon}$ with the new candidate $G_{s}$ only when this leads to a $(1+\frac{1}{k})$ factor improvement for $\sf DS$}. This ensures a $\frac{1}{4}$-approximation ratio [49]. Following the replacement policy consistently,  updateSX ensures a $\frac{1}{4}$ approximation ratio for $(\zeta^{\prime},\epsilon)$-explanations in $\zeta^{\prime}$. $\Box$
Time Cost. As Algorithm $\mathsf{ASX}$-$\mathsf{OP}$ processes candidate subgraphs in an online fashion from a stream, we analyze its time complexity in an output-sensitive manner in terms of $\zeta^{\prime}$ at termination. It evaluates $|\zeta^{\prime}|$ candidate subgraphs. For each candidate $G_{s}$, the $\mathsf{Verifier}$ module (procedures vrfyF and vrfyCF) checks its validity for $\mathsf{SXQ}$ via two forward passes of the $\mathsf{GNN}$ model $\mathcal{M}$, incurring a cost of $O(L|G^{L}(v_{t})|)$ per candidate, assuming a small number of node features (cf. Lemma III-C[43, 37]). Hence, the total verification cost is $O(|\zeta^{\prime}|L|G^{L}(v_{t})|)$. Verified candidates are passed to the $\mathsf{Updater}$ module ($\sf updateSX$}), which maintains a $(1+\epsilon)$-approximate Pareto set in $\zeta^{\prime}$ by solving a multi-objective dominance update over the state space. This takes at most $O\left(\prod_{i=1}^{|\Phi|}\left(\left\lfloor\log_{1+\epsilon}\frac{\phi_{u}}{\phi_{l}}\right\rfloor+1\right)\right)$ time per update, where $\phi_{l}$ and $\phi_{u}$ are the minimum and maximum values observed for each metric $\phi\in\Phi$. Approximating $\log_{1+\epsilon}\approx\frac{1}{\epsilon}$ for small $\epsilon$, the total maintenance cost becomes $O\left(|\zeta^{\prime}|\cdot\left(\frac{\log r_{\Phi}}{\epsilon}\right)^{|\Phi|}\right)$, where $r_{\Phi}=\frac{\phi_{u}}{\phi_{l}}$. Therefore, the total time cost of $\mathsf{ASX}$-$\mathsf{OP}$ is in $O\left(|\zeta^{\prime}|\cdot\left(\left(\frac{\log r_{\Phi}}{\epsilon}\right)^{|\Phi|}+L|G^{L}(v_{t})|\right)\right)$.
@@ IV-B Diversified Skyline Explanations
Skyline explanation may still contain explanatory subgraphs having highly “similar” or overlapped nodes. This may lead to redundant and biased explanations. In practice, diversified explanations often clarify the model output more comprehensively [50, 51]. We next investigate a diversified scheme for ${\mathsf{SXQ}}^{k}$ evaluation, in terms of coverage and the difference between node representations.
@@
where $\mathbf{x}_{G_{s}}$ is the embedding of $G_{s}$ obtained by node embedding learning such as Node2Vec [52]. The two terms are balanced by a constant $\alpha$.
Diversification Algorithm. We next outline our diversified evaluation algorithm, denoted as $\mathsf{DSX}$ (pseudo-code shown in [47]). It follows $\mathsf{ASX}$-$\mathsf{OP}$ and adopts onion peeling strategy to generate subgraphs in a stream. The difference is that when computing the $k$-skyline explanation, it includes a new replacement strategy when the marginal gain for the diversification function ${\sf DivS}(\cdot)$ is at least a factor of the score over the current solution ${\sf DivS}({\mathcal{G}}_{\epsilon})$. $\mathsf{DSX}$ terminates when a first $k$-skyline explanation is found.
@@
We show the above result by reducing the diversified evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the Streaming Submodular Maximization problem [53]. The problem maintains a size-$k$ set that optimizes a submodular function over a stream of data objects. We show that $\mathsf{DSX}$ adopts a consistent increment policy that ensures a ($\frac{1}{2}-\epsilon$)-approximation as in [53]. We present the detailed proof in [47].
@@ V Parallel Skyline Explanatory Querying
Parallel Algorithm. The parallel algorithm, denoted as $\mathsf{ParaSX}$, works with $m$ threads to evaluate a query set $Q$ with $n$ $\mathsf{SXQ}$} queries, involving a test node set $V_{T}$ (pseudocode given in the full version [47]). (1) Inter-Batch parallelization: It first partitions $Q$ into $N$ small batches to be evaluated in parallel, guided by a clustering strategy with a goal that for any pair of target nodes $v_{t}$ and $v_{t}^{\prime}$ from two different batches, their $L$-hop neighbors have minimized overlap (see “Workload Clustering”). (2) Batch Assignment. $\mathsf{ParaSX}$ then solves a makespan minimization problem to assign the $N$ batches to $m$ threads, with a dynamic scheduling strategy that prioritizes the processing of queries with “influencing” output nodes that have highest common $L$-hop neighbors for those in the same thread. This is to maximize the shared computation for edge peeling and verification. (3) Inter-query Pipelining: Each local thread then invokes $\mathsf{ASX}$-$\mathsf{OP}$ ($v_{t}$) to generate local $k$-explanations for each assigned node $v_{t}$. All the processing follows its local ranks of highly influential nodes, and automatically skips processing a node, if its information has been updated via peeling operations earlier by previous nodes. This processing strategy is enabled by maintaining a shared data structure that records nodes and their $L$-hop neighboring edges within each thread. The skyline computation can be further parallelized by parallel skyline query methods [54, 55].
@@
$\mathsf{Cora}$ [56] 2,708 10,556 1,433 7
$\mathsf{PubMed}$ [57] 19,717 88,648 500 3
$\mathsf{FacebookPage}$ [58] 22,470 342,004 128 4
$\mathsf{AmazonComputer}$ [59] 13,752 491,722 767 10
$\mathsf{OGBN\_arxiv}$ [60] 169,343 1,166,243 128 40
$\mathsf{BA\_Shapes}$ [6] 1,010,000,000 6,027,852,000 1 4
$\mathsf{Tree\_Cycles}$ [6] 991 2140 1 2

### v4

@@ III-A Skyline Explanatory Query
That is, ${\mathcal{G}}_{\zeta}$ is a Pareto set [sharma2022comprehensive] of the interpretable space $\zeta$.
@@ III-C Computational Complexity
Proof sketch: Given an $\mathsf{SXQ}$} = $(G,{\mathcal{M}},v_{t},\Phi)$ and a set of subgraphs $\mathcal{G}$} of $G$, we provide a procedure, denoted as $\mathsf{Verify}$, that correctly determines if $\mathcal{G}$} is an explanation. The algorithm checks, for each $G_{s}\in{\mathcal{G}}$, if $G_{s}$ is a factual or a counterfactual explanation of ${\mathcal{M}}(v_{t},G)$. It has been verified that this process can be performed by invoking a polynomial time inference process of $\mathcal{M}$} [chen2020scalable, robogexp] for $v_{t}$ over $G_{s}$ (for testing factual explanation) and $G\setminus G_{s}$ (for testing counterfactual explanations), respectively. $\Box$
@@
Proof sketch: It suffices to show that ${\mathsf{EVAL}}({\mathsf{SXQ}}^{k})$ is $\mathsf{NP}$-hard for an interpretable space $|\zeta|$ = $f(|G|)$ for some polynomial function $f$, and $|\Phi|=3$. We show this by constructing a polynomial-time reduction from $k$-representative skyline selection problem ($k$-RSP). Given a set $S$ of $d$-dimensional data points, $k$-RSP is to compute a subset $S^{*}$ of $S$, such that (a) $S^{*}$ is a set of skyline points, $|S^{*}|=k$, and (b) $S^{*}$ maximizes the dominance score. $k$-RSP is $\mathsf{NP}$-hard even for $3$-dimensional data points [lin2006selecting].
@@
A straightforward evaluation of ${\mathsf{SXQ}}^{k}$ may explore the $L$-hop subgraph $G^{L}(\{v_{t}\})$ and initializes the interpretable space $\zeta$ as all connected subgraphs in $G^{L}(\{v_{t}\})$, by invoking subgraph enumeration algorithms [karakashian2013algorithm, Yang21]. It then enumerates $n$ size-$k$ subsets of $\zeta$ and finds an optimal $k$-explanation. Although this correctly finds optimal explanations, it is not practical for large $G$, as $n$ alone can be $2^{deg^{L}}$ and the Pareto sets inspection can take $O(\binom{n}{k})$ time. We thus consider “approximate query processing” scheme for ${\mathsf{SXQ}}^{k}$, and present efficient algorithms that do not require such enumeration.
@@ IV-A Approximating Skyline Explanations
In other words, a $(\zeta,\epsilon)$-explanation ${\mathcal{G}}_{\epsilon}$ approximates a $k$-explanation ${\mathcal{G}}_{\zeta}$ as its answer in the interpretable space $\zeta$. Indeed, (1) ${\mathcal{G}}_{\epsilon}$ has a bounded number $k$ of explanatory subgraphs as ${\mathcal{G}}_{\zeta}$; (2) ${\mathcal{G}}_{\epsilon}$ is, by definition, an $\epsilon$-Pareto set [schutze2021computing] of $\zeta$. In multi-objective decision making, $\epsilon$-Pareto sets have been verified as a class of cost-effective, size-bounded approximation for Pareto optimal solutions.
@@
Update Dominance Relations. Procedure updateSX maintains a set of skyline explanations ${\mathcal{G}}_{s}$ (not shown) in terms of $(1+\epsilon)$-dominance. The set ${\mathcal{G}}_{s}$ is efficiently derived from the bitvectors $B(s)$ of the states in $\zeta^{\prime}$. The latter compactly encodes a lattice structure of $(1+\epsilon)$ dominance as a directed acyclic graph; and ${\mathcal{G}}_{s}$ refers to the states with no “parent” in the lattice, i.e., having an explanatory subgraph that is not $(1+\epsilon)$-dominated by any others so far. Details are in Appendix.
@@
Proof sketch: We show the above result with a reduction to the multi-objective shortest path problem ($\mathsf{MOSP}$[tsaggouris2009multiobjective]. Given an edge-weighted graph $G_{w}$, where each edge carries a $d$-dimensional attribute vector $e_{w}.c$, it computes a Pareto set of paths from a start node $u$. The cost of a path $\rho_{w}$ in $G_{w}$ is defined as $\rho_{w}.c$ = $\sum_{e_{w}\in\rho_{w}}$ $e_{w}.c$. The dominance relation between two paths is determined by the dominance relation of their cost vector. Our reduction (1) constructs $G_{w}$ as the running graph $\zeta^{\prime}$ with $n$ verified states and transitions; and (2) for each edge $(s,s^{\prime})$, sets an edge weight as $e_{w}$ = $\Phi(s)-\Phi(s^{\prime})$. Given a solution $\Pi_{w}$ of the above instance of $\mathsf{MOSP}$, for each path $\rho_{w}\in\Pi$, we set a corresponding path $\rho$ in $\zeta^{\prime}$ that ends at a state $\rho_{s}$ , and adds it into ${\mathcal{G}}_{\epsilon}$. We can verify that $\Pi_{w}$ is an $\epsilon$-Pareto set of paths $\Pi_{w}$ in $G_{w}$, if and only if ${\mathcal{G}}_{\epsilon}$ is an $(\zeta^{\prime},\epsilon)$-explanation of $\zeta^{\prime}$. We show that $\mathsf{ASX}$-$\mathsf{OP}$ performs a simpler process of the algorithm in [tsaggouris2009multiobjective], which ensures to generate ${\mathcal{G}}_{\epsilon}$ as a $(\zeta^{\prime},\epsilon)$-explanation. $\Box$
@@
Proof sketch: Consider procedure updateSX upon the arrival, at any time, of a new verified candidate $G_{s}$. (1) The above results clearly hold when $|\zeta^{\prime}|\leq k$ or $|{\mathcal{G}}_{\epsilon}|\leq k$, as ${\mathcal{G}}_{\epsilon}$ is the only $(\zeta^{\prime},\epsilon)$-explanation so far. (2) When $|{\mathcal{G}}_{\epsilon}|=k$, we reduce the approximate evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the online MAX $k$-SET problem [onlinemkc]. The problem maintains a size-$k$ set cover with maximized weights. We show that updateSX adopts a greedy replacement policy by replacing a candidate in ${\mathcal{G}}_{\epsilon}$ with the new candidate $G_{s}$ only when this leads to a $(1+\frac{1}{k})$ factor improvement for $\sf DS$}. This ensures a $\frac{1}{4}$-approximation ratio [onlinemkc]. Following the replacement policy consistently,  updateSX ensures a $\frac{1}{4}$ approximation ratio for $(\zeta^{\prime},\epsilon)$-explanations in $\zeta^{\prime}$. $\Box$
Time Cost. As Algorithm $\mathsf{ASX}$-$\mathsf{OP}$ processes candidate subgraphs in an online fashion from a stream, we analyze its time complexity in an output-sensitive manner in terms of $\zeta^{\prime}$ at termination. It evaluates $|\zeta^{\prime}|$ candidate subgraphs. For each candidate $G_{s}$, the $\mathsf{Verifier}$ module (procedures vrfyF and vrfyCF) checks its validity for $\mathsf{SXQ}$ via two forward passes of the $\mathsf{GNN}$ model $\mathcal{M}$, incurring a cost of $O(L|G^{L}(v_{t})|)$ per candidate, assuming a small number of node features (cf. Lemma III-C[chen2020scalable, robogexp]). Hence, the total verification cost is $O(|\zeta^{\prime}|L|G^{L}(v_{t})|)$. Verified candidates are passed to the $\mathsf{Updater}$ module ($\sf updateSX$}), which maintains a $(1+\epsilon)$-approximate Pareto set in $\zeta^{\prime}$ by solving a multi-objective dominance update over the state space. This takes at most $O\left(\prod_{i=1}^{|\Phi|}\left(\left\lfloor\log_{1+\epsilon}\frac{\phi_{u}}{\phi_{l}}\right\rfloor+1\right)\right)$ time per update, where $\phi_{l}$ and $\phi_{u}$ are the minimum and maximum values observed for each metric $\phi\in\Phi$. Approximating $\log_{1+\epsilon}\approx\frac{1}{\epsilon}$ for small $\epsilon$, the total maintenance cost becomes $O\left(|\zeta^{\prime}|\cdot\left(\frac{\log r_{\Phi}}{\epsilon}\right)^{|\Phi|}\right)$, where $r_{\Phi}=\frac{\phi_{u}}{\phi_{l}}$. Therefore, the total time cost of $\mathsf{ASX}$-$\mathsf{OP}$ is in $O\left(|\zeta^{\prime}|\cdot\left(\left(\frac{\log r_{\Phi}}{\epsilon}\right)^{|\Phi|}+L|G^{L}(v_{t})|\right)\right)$.
@@ IV-B Diversified Skyline Explanations
Skyline explanation may still contain explanatory subgraphs having highly “similar” or overlapped nodes. This may lead to redundant and biased explanations. In practice, diversified explanations often clarify the model output more comprehensively [gong2019diversity, mothilal2020explaining]. We next investigate a diversified scheme for ${\mathsf{SXQ}}^{k}$ evaluation, in terms of coverage and the difference between node representations.
@@
where $\mathbf{x}_{G_{s}}$ is the embedding of $G_{s}$ obtained by node embedding learning such as Node2Vec [grover2016node2vec]. The two terms are balanced by a constant $\alpha$.
Diversification Algorithm. We next outline our diversified evaluation algorithm, denoted as $\mathsf{DSX}$ (pseudo-code shown in Appendix). It follows $\mathsf{ASX}$-$\mathsf{OP}$ and adopts onion peeling strategy to generate subgraphs in a stream. The difference is that when computing the $k$-skyline explanation, it includes a new replacement strategy when the marginal gain for the diversification function ${\sf DivS}(\cdot)$ is at least a factor of the score over the current solution ${\sf DivS}({\mathcal{G}}_{\epsilon})$. $\mathsf{DSX}$ terminates when a first $k$-skyline explanation is found.
@@
We show the above result by reducing the diversified evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the Streaming Submodular Maximization problem [ssm]. The problem maintains a size-$k$ set that optimizes a submodular function over a stream of data objects. We show that $\mathsf{DSX}$ adopts a consistent increment policy that ensures a ($\frac{1}{2}-\epsilon$)-approximation as in [ssm]. We present the detailed proof in Appendix.
@@ V Parallel Skyline Explanatory Querying
Parallel Algorithm. The parallel algorithm, denoted as $\mathsf{ParaSX}$, works with $m$ threads to evaluate a query set $Q$ with $n$ $\mathsf{SXQ}$} queries, involving a test node set $V_{T}$ (pseudocode given in Appendix). (1) Inter-Batch parallelization: It first partitions $Q$ into $N$ small batches to be evaluated in parallel, guided by a clustering strategy with a goal that for any pair of target nodes $v_{t}$ and $v_{t}^{\prime}$ from two different batches, their $L$-hop neighbors have minimized overlap (see “Workload Clustering”). (2) Batch Assignment. $\mathsf{ParaSX}$ then solves a makespan minimization problem to assign the $N$ batches to $m$ threads, with a dynamic scheduling strategy that prioritizes the processing of queries with “influencing” output nodes that have highest common $L$-hop neighbors for those in the same thread. This is to maximize the shared computation for edge peeling and verification. (3) Inter-query Pipelining: Each local thread then invokes $\mathsf{ASX}$-$\mathsf{OP}$ ($v_{t}$) to generate local $k$-explanations for each assigned node $v_{t}$. All the processing follows its local ranks of highly influential nodes, and automatically skips processing a node, if its information has been updated via peeling operations earlier by previous nodes. This processing strategy is enabled by maintaining a shared data structure that records nodes and their $L$-hop neighboring edges within each thread. The skyline computation can be further parallelized by parallel skyline query methods [wu2006parallelizing, park2013parallel].
@@
$\mathsf{Cora}$ [cora] 2,708 10,556 1,433 7
$\mathsf{PubMed}$ [pubmed] 19,717 88,648 500 3
$\mathsf{FacebookPage}$ [facebook] 22,470 342,004 128 4
$\mathsf{AmazonComputer}$ [shchur2018pitfalls] 13,752 491,722 767 10
$\mathsf{OGBN\_arxiv}$ [hu2020open] 169,343 1,166,243 128 40
$\mathsf{BA\_Shapes}$ [ying2019gnnexplainer] 1,010,000,000 6,027,852,000 1 4
$\mathsf{Tree\_Cycles}$ [ying2019gnnexplainer] 991 2140 1 2